OpenStudy (anonymous):

A volcano launches a lava bomb straight upward with an initial speed of 28m/s. Taking upward to be the positive direction,find speed and direction of motion of the lava bomb a) 2.0seconds and b)3.0seconds after it is launched

OpenStudy (anonymous):

v = u +gt a) u = 28 m/s g = 10 m/s^2 (actually it is 9.8 but for ease of calculation we take it as 10) t = 2s v = 28 m/s + 10 m/s^2 * 2s = 28 m/s + 20 m/s = 48 m/s after 2s the lava bomb is travelling upwards with a speed of 48 m/s b) u = 28 m/s g = 10 m/s^2 (actually it is 9.8 but for ease of calculation we take it as 10) t = 3s v = 28 m/s + 10 m/s^2 * 3s = 28 m/s + 30 m/s = 58 m/s after 3s the lave bomb is travelling upwards with a speed of 58 m/s

OpenStudy (anonymous):

You have to make g negative, since it's acting towards the earth. a)v = 28 -10*2 = 8m/s (still going up, put with slower speed) b)v = 28 -10*3 =-2m/s (coming down, picking up speed)

OpenStudy (anonymous):

yep